The assembly steps/times to create a single bicycle sums to 195 seconds. The production time per day is 25,000 seconds and output per day is 600 bicycles.

If, after completing an assembly re-design, you require 6 stations, what is the percent efficiency? (note, you only need to enter the number)
Answer #1

Please note :

Required cycle time = Available time per day / Required production quantity per day

Efficiency of assembly line = (Total time required at all steps for a product to come out / Actual number of station x Cycle time ) x 100

Accordingly,

Cycle time = 25,000/600 = 41.67 ( rounded to 2 decimal places )

Efficiency of assembly line =   195 / ( 6 x 41.67) x 100 = 77.993 %

PERCENTAGE EFFICIENCY = 77.993%